OpenStudy (anonymous):

MEDAL & FAN! what is the volume of a cube If each edge of that cube is a length of 1/4 of a centimeter?

OpenStudy (johnweldon1993):

Hint* volume = length x width x height For a cube...all those values are equal... So \[\large V = s^3 = (\frac{1}{4}cm)^3\]

OpenStudy (er.mohd.amir):

side =1/4 of cm =.25 cm volume of cube = side*side * side

OpenStudy (anonymous):

.25 cm is the answer?

OpenStudy (er.mohd.amir):

no vol=0.25*0.25*0.25 cm^3

OpenStudy (anonymous):

0.015625

OpenStudy (er.mohd.amir):

yes